OpenStudy (anonymous):

Consider the set of all points less than 5 units away from the origin. a. Write an inequality to represent the points.

OpenStudy (skullpatrol):

Any ideas?

ganeshie8 (ganeshie8):



OpenStudy (anonymous):

^ I was thinking that, but on 4?

ganeshie8 (ganeshie8):

what do u mean 4

OpenStudy (anonymous):

4 points away, since it says less than 5

ganeshie8 (ganeshie8):

good thinking, but no

ganeshie8 (ganeshie8):

here we need to write an inequality

ganeshie8 (ganeshie8):

points that are 5 units away from center form a circle wid radius 5 units, eh ?

OpenStudy (anonymous):

mhm

ganeshie8 (ganeshie8):

whats the equation of circle with centern(0, 0) and radius 5 ?

OpenStudy (anonymous):

\[(x-0)^2 + (y-0)^2 = 5\]?

ganeshie8 (ganeshie8):

put 5^2

ganeshie8 (ganeshie8):

\(x^2 + y^2 = 5^2 \)

ganeshie8 (ganeshie8):